OpenStudy (anonymous):

Consider the arc length of the curve y = ln x on 1<= x <= 2. a) First, set up, but do not evaluate, an integral that represents the exact arc length. b) Next, apply the u-substitution, u = 1/x, to your integral in part (a). Leave your integral in terms of u only without evaluating the integral. c) Now, apply the trigonometric substitution u = tan(theta) to your integral in part (b), and find the arc length.

ganeshie8 (ganeshie8):

arc length = \(\large \mathbb{\int_1^2 \sqrt{1 + \left(\frac{dy}{dx}\right)^2} dx}\)

ganeshie8 (ganeshie8):

find the derivative and plugin

OpenStudy (anonymous):

I get \[\int\limits_{1}^{2} \frac{ 1 }{ x } \sqrt{x^2+1}dx\] for (a). Im suppose to u-sub u=1/x for (b) but I get confused here

ganeshie8 (ganeshie8):

just do watever it is asking to do.. u not going to evaluate it anyways

OpenStudy (anonymous):

I get u = 1/x \[du = \frac{ -1 }{ x^2 }\] I don't know where to plug in the du

ganeshie8 (ganeshie8):

sub \( u = \frac{1}{x}\) \(du = \frac{-1}{x^2} dx\) \(-xdu = \frac{1}{x} dx\) \(-\frac{1}{u}du = \frac{1}{x} dx\)

ganeshie8 (ganeshie8):

as x->1, u->1 as x->2, u->1/2

ganeshie8 (ganeshie8):

so the integral becomes : \(\large \int\limits_{1}^{\frac{1}{2}} \sqrt{(\frac{1}{u})^2+1} (\frac{-du}{u}) \)

ganeshie8 (ganeshie8):

leave it like that...

OpenStudy (anonymous):

Then I have to use trig sub

OpenStudy (anonymous):

u = tan(theta)

ganeshie8 (ganeshie8):

ya go ahead